OpenStudy (anonymous):

2x-17=9/x solve the equation by factoring

jimthompson5910 (jim_thompson5910):

2x-17=9/x x(2x-17)=9 2x^2 - 17x = 9 2x^2 - 17x - 9 = 0 2x^2+x-18x-9 = 0... See note below (2x^2+x)+(-18x-9) = 0 x(2x+1)+(-18x-9) = 0 x(2x+1)-9(2x+1) = 0 (x-9)(2x+1) = 0 x-9=0 or 2x+1=0 I'll let you finish

jimthompson5910 (jim_thompson5910):

note: you multiply 2 and -9 to get -18 then you find two numbers that multiply to -18 AND add to -17. These two numbers are 1 and -18, so that's how I broke -17x into x - 18x